OpenStudy (anonymous):

A circle is represented by the equation shown below. (x + 3)^2 + (y - 2)^2 = 4 Which statement is true? The circle is centered at (3, 2) and has a radius of 2. The circle is centered at (-3, 2) and has a radius of 2. The circle is centered at (-3, -2) and has a diameter of 2. The circle is centered at (-3, 2) and has a diameter of 2.

OpenStudy (anonymous):

i know its B or D

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

=r^2

OpenStudy (anonymous):

so is it the radius or diameter that is 2?

OpenStudy (anonymous):

if it's =4, then you have 2^2 since I said the equation is r^2 then you have 2^2=r^2

OpenStudy (campbell_st):

ok.. so the general form is \[(x -h)^2 + (y - k)^2 = r^2\] the centre is (h, k) and the radius is r

OpenStudy (campbell_st):

just match your question to the general form

OpenStudy (anonymous):

so the answer is "The circle is centered at (-3, 2) and has a radius of 2."

OpenStudy (campbell_st):

thats correct.

OpenStudy (anonymous):

the answer is B